# for Human Motion Generation

### v2

@@ Abstract.
Human motion generation has found widespread applications in AR/VR, film, sports, and medical rehabilitation, offering a cost-effective alternative to traditional motion capture systems. However, evaluating the fidelity of such generated motions is a crucial, multifaceted task. Although previous approaches have attempted at motion fidelity evaluation using human perception or physical constraints, there remains an inherent gap between human-perceived fidelity and physical feasibility. Moreover, the subjective and coarse binary labeling of human perception further undermines the development of a robust data-driven metric. We address these issues by introducing a physical labeling method. This method evaluates motion fidelity by calculating the minimum modifications needed for a motion to align with physical laws. With this approach, we are able to produce fine-grained, continuous physical alignment annotations that serve as objective ground truth. With these annotations, we propose $P$-Motion, a novel data-driven metric to evaluate both physical and perceptual fidelity of human motion. To effectively capture underlying physical priors, we employ Pearson’s correlation loss for the training of our metric. Additionally, by incorporating a human-based perceptual fidelity loss, our metric can capture fidelity that simultaneously considers both human perception and physical alignment. Experimental results demonstrate that our metric, $P$-Motion, not only aligns with physical laws but also aligns better with human perception of motion fidelity than previous work. Project page: https://sarah816.github.io/$p$-motion-site/.
@@
Figure 1. A motion that looks realistic does not necessarily mean it is physically feasible. Top-left: Motion appears realistic and semantically meaningful to the human eye, yet fails in physics simulation, resulting in a fall (bottom-left). Top-right: Unnatural motion in human perception executes successfully in simulation (bottom-right). This reveals a discrepancy between human perception and physical laws.:
@@ 1. Introduction
To address these two issues, we propose a data-driven method to evaluate motion fidelity that aligns with physical laws. We achieve this by calculating the minimum distance between the test motion and a motion that complies with physical laws. A small minimal distance indicates a high fidelity of the input motion, whereas a large minimal distance implies low fidelity. With this physically grounded definition of fidelity, we can establish fine-grained continuous labels for physical law alignments. In addition, we design a framework that trains a metric, named $P$-Motion, by simultaneously utilizing fine-grained physical labels and coarse, discrete human perceptual labels. By designing loss functions that better suit the fine-grained labels, we can more effectively learn the underlying physical law priors. With this approach, we are able to train our metric to better align with physical laws. Furthermore, since human perception is inherently correlated with physical feasibility, this approach also improves the potential for metric design to align with human judgments by learning physical principles.
@@
  • We propose a novel fidelity evaluation method, $P$-Motion, for human motions, which takes into account both physical feasibility and human perception. Our method can evaluate whether a motion is realistically aligned with physical laws and human perception.
@@
Figure 2. Our metric design and training pipeline. The network takes a human motion sequence as input, which is processed by a motion encoder to extract spatiotemporal features. These features are then decoded into a fidelity score by a fidelity decoder. The network is trained in a supervised fashion using fine-grained physical annotations alongside human perceptual labels. :
@@ 2.1. Human Motion Generation
$2.$ Human Motion Evaluation
@@ 3.2. Physical-Perceptual Motion Metric
Motion encoder. The motion encoder plays a critical role in determining the quality of the motion features. To extract both spatial and temporal information necessary for fidelity evaluation, we adopt a state-of-the-art spatio-temporal motion encoder. In our experiments, we follow the design proposed in (Zhu et al., 2023a). This encoder is built from NN dual-stream fusion modules, each containing branches for spatial and temporal self-attention and MLP. The spatial layers capture correlations among different joints within the same time step, while the temporal layers focus on the dynamics of individual joints. This dual-stream design effectively captures comprehensive features required for assessing motion fidelity.
@@
Our approach to physical supervision primarily focuses on aligning the network’s output with the physical annotations. We have observed that, for designing a physically aligned fidelity score, the absolute value of the output score is less critical than its correlation with the physical labels. Therefore, our supervision targets the correlation between the fidelity score and the physical annotations rather than direct numerical differences as in traditional regression tasks. To this end, we use Pearson’s correlation loss as the core training loss, as detailed in Sec. $3.$
@@
Figure 3. Visualized examples for our dataset labeling. Above: Original motion sequence from the MotionPercept dataset. Below: The motion sequence imitated in the simulator using a per-sequence fine-tuned model. :
@@ 4.2. Physical Annotation Generation
Figure 4. (a) A MotionPercept better/worse data pair: the motion on the left, annotated as ‘better’ and visually superior, exhibits physics issues (e.g. floating, skating). The motion on the right, annotated as ‘worse’ and visually inferior, shows greater physical plausibility. (b) Three MotionPercept samples in a group, all annotated as ‘worse’ by humans, reveal different physical characteristics in the simulator. Our $P$-Motion scores (higher means better) successfully capture these physical distinctions. :
@@ 5. Experiment Results
Comparison with previous metrics. We verify the human perceptual and physical alignment of our metric $P$-Motion and previous metrics on 2 subsets of MotionPercept: MDM and FLAME. The results are shown in Tab. 3. For human perceptual alignment, we evaluated the “better/worse” classification accuracy. For physical alignment, we evaluated 3 correlation methods (i.e., PLCC (Pearson, 1920), SROCC (Spearman, 1910), KROCC (Loshchilov and Hutter, 2017) between $P$-Motion results and physical annotation. The correlation methods are defined in supplementary materials. Our metric is trained only on the MDM training set and directly tested on the MotionPercept-MDM validation and the MotionPercept-FLAME dataset without further finetuning. The results show that our metric outperforms previous works on physical alignment. Notably, on human perceptual alignment, our metric slightly outperforms the baseline method, MotionCritic, which proves that the physical annotation has the potential to improve the human perceptual alignment of our metric. We also compare our metric with existing pose-based metrics (i.e., Root AVE, Root AE, Joint AVE, Joint AE) and physics-based metrics (i.e., Penetration, Float, Skate). Detailed definitions of these metrics are provided in supplementary materials.
@@
Improving motion generation with $P$-Motion. We try to improve the motion generation method MDM (Tevet et al., 2022b) to further verify our metric’s physical fidelity alignment. We fine-tune the MDM network using our metric, along with critic loss and KL loss from (Wang et al., 2025). The critic loss for the input pose $x^{\prime}_{0}$ is defined as:
@@
Visualization. Fig. 4 (a) shows a better/worse data pair sampled from the MotionPercept dataset. The motion on the left (annotated as ‘better’ and visually superior) exhibits physics issues (e.g. floating, skating) in the simulator, while the motion on the right (annotated as ‘worse’ and visually inferior) demonstrates greater physical plausibility when simulated. Fig. 4 (b) shows three MotionPercept samples in a group that are all annotated as ‘worse’, but reveal different physical characteristics in the simulator. Our $P$-Motion scores successfully capture these physical distinctions. Note that for our $P$-Motion score, the higher reveals the better.
@@ Appendix A Implementation Details
We train our $P$-Motion model using the MDM subset of MotionPercept, which contains 46761 better-worse motion pairs. The MDM dataset is generated from 12 action labels in HumanAct12 and 40 labels in UESTC. We first categorize the dataset according to 52 prompt labels. During training, each batch contains motions exclusively from the same category, and the PLCC loss is computed collectively across all motions within a batch. Our model architecture follows (Wang et al., 2025), employing a DSTformer (Zhu et al., 2023b) backbone with 3 layers and 8 attention heads for generating motion embedding. The embeddings are then fed into a 1024-channel MLP layer that produces a single scalar score. We train the $P$-Motion model for 200 epochs with a batch size of 64 and a learning rate initialized at 4e-5 and decayed exponentially (factor 0.995 per epoch). The correlation loss term is weighted by $\lambda=0.3$.
@@ D.2. Visualize cases
We provide visualized cases from the MotionPercept dataset in Fig. 4 to verify the effectiveness of our $P$-Motion metric.
@@
Our $P$-Motion scoring effectively captures these physical plausibility considerations:
@@
$P$-Motion scoring demonstrates strong correlation with physical feasibility assessments:

### v3

# PP-Motion: Physical-Perceptual Fidelity Evaluation for Human Motion Generation

@@ Abstract.
Human motion generation has found widespread applications in AR/VR, film, sports, and medical rehabilitation, offering a cost-effective alternative to traditional motion capture systems. However, evaluating the fidelity of such generated motions is a crucial, multifaceted task. Although previous approaches have attempted at motion fidelity evaluation using human perception or physical constraints, there remains an inherent gap between human-perceived fidelity and physical feasibility. Moreover, the subjective and coarse binary labeling of human perception further undermines the development of a robust data-driven metric. We address these issues by introducing a physical labeling method. This method evaluates motion fidelity by calculating the minimum modifications needed for a motion to align with physical laws. With this approach, we are able to produce fine-grained, continuous physical alignment annotations that serve as objective ground truth. With these annotations, we propose PP-Motion, a novel data-driven metric to evaluate both physical and perceptual fidelity of human motion. To effectively capture underlying physical priors, we employ Pearson’s correlation loss for the training of our metric. Additionally, by incorporating a human-based perceptual fidelity loss, our metric can capture fidelity that simultaneously considers both human perception and physical alignment. Experimental results demonstrate that our metric, PP-Motion, not only aligns with physical laws but also aligns better with human perception of motion fidelity than previous work. Code is available at: https://github.com/Sarah816/PP-Motion.
@@
Figure 1. A motion that looks realistic does not necessarily mean it is physically feasible. Top-left: Motion appears realistic and semantically meaningful to the human eye, yet fails in physics simulation, resulting in a fall (bottom-left). Top-right: Unnatural motion in human perception executes successfully in simulation (bottom-right). This reveals a discrepancy between human perception and physical laws.
A teaser figure to show our motivation.
@@ 1. Introduction
To address these two issues, we propose a data-driven method to evaluate motion fidelity that aligns with physical laws. We achieve this by calculating the minimum distance between the test motion and a motion that complies with physical laws. A small minimal distance indicates a high fidelity of the input motion, whereas a large minimal distance implies low fidelity. With this physically grounded definition of fidelity, we can establish fine-grained continuous labels for physical law alignments. In addition, we design a framework that trains a metric, named PP-Motion, by simultaneously utilizing fine-grained physical labels and coarse, discrete human perceptual labels. By designing loss functions that better suit the fine-grained labels, we can more effectively learn the underlying physical law priors. With this approach, we are able to train our metric to better align with physical laws. Furthermore, since human perception is inherently correlated with physical feasibility, this approach also improves the potential for metric design to align with human judgments by learning physical principles.
@@
  • We propose a novel fidelity evaluation method, PP-Motion, for human motions, which takes into account both physical feasibility and human perception. Our method can evaluate whether a motion is realistically aligned with physical laws and human perception.
@@
Figure 2. Our metric design and training pipeline. The network takes a human motion sequence as input, which is processed by a motion encoder to extract spatiotemporal features. These features are then decoded into a fidelity score by a fidelity decoder. The network is trained in a supervised fashion using fine-grained physical annotations alongside human perceptual labels.
Pipeline
@@ 2.1. Human Motion Generation
2.2. Human Motion Evaluation
@@ 3.2. Physical-Perceptual Motion Metric
Motion encoder. The motion encoder plays a critical role in determining the quality of the motion features. To extract both spatial and temporal information necessary for fidelity evaluation, we adopt a state-of-the-art spatio-temporal motion encoder. In our experiments, we follow the design proposed in (Zhu et al., 2023a). This encoder is built from $N$ dual-stream fusion modules, each containing branches for spatial and temporal self-attention and MLP. The spatial layers capture correlations among different joints within the same time step, while the temporal layers focus on the dynamics of individual joints. This dual-stream design effectively captures comprehensive features required for assessing motion fidelity.
@@
Our approach to physical supervision primarily focuses on aligning the network’s output with the physical annotations. We have observed that, for designing a physically aligned fidelity score, the absolute value of the output score is less critical than its correlation with the physical labels. Therefore, our supervision targets the correlation between the fidelity score and the physical annotations rather than direct numerical differences as in traditional regression tasks. To this end, we use Pearson’s correlation loss as the core training loss, as detailed in Sec. 3.3.
@@ 3.3. Training Loss
Figure 3. Visualized examples for our dataset labeling. Above: Original motion sequence from the MotionPercept dataset. Below: The motion sequence imitated in the simulator using a per-sequence fine-tuned model.
Visualized examples for our dataset labeling.
@@ 4.2. Physical Annotation Generation
Figure 4. (a) A MotionPercept better/worse data pair: the motion on the left, annotated as ‘better’ and visually superior, exhibits physics issues (e.g. floating, skating). The motion on the right, annotated as ‘worse’ and visually inferior, shows greater physical plausibility. (b) Three MotionPercept samples in a group, all annotated as ‘worse’ by humans, reveal different physical characteristics in the simulator. Our PP-Motion scores (higher means better) successfully capture these physical distinctions.
Visualization
@@ 5. Experiment Results
Comparison with previous metrics. We verify the human perceptual and physical alignment of our metric PP-Motion and previous metrics on 2 subsets of MotionPercept: MDM and FLAME. The results are shown in Tab. 3. For human perceptual alignment, we evaluated the “better/worse” classification accuracy. For physical alignment, we evaluated 3 correlation methods (i.e., PLCC (Pearson, 1920), SROCC (Spearman, 1910), KROCC (Loshchilov and Hutter, 2017) between PP-Motion results and physical annotation. The correlation methods are defined in supplementary materials. Our metric is trained only on the MDM training set and directly tested on the MotionPercept-MDM validation and the MotionPercept-FLAME dataset without further finetuning. The results show that our metric outperforms previous works on physical alignment. Notably, on human perceptual alignment, our metric slightly outperforms the baseline method, MotionCritic, which proves that the physical annotation has the potential to improve the human perceptual alignment of our metric. We also compare our metric with existing pose-based metrics (i.e., Root AVE, Root AE, Joint AVE, Joint AE) and physics-based metrics (i.e., Penetration, Float, Skate). Detailed definitions of these metrics are provided in supplementary materials.
@@
Improving motion generation with PP-Motion. We try to improve the motion generation method MDM (Tevet et al., 2022b) to further verify our metric’s physical fidelity alignment. We fine-tune the MDM network using our metric, along with critic loss and KL loss from (Wang et al., 2025). The critic loss for the input pose $x^{\prime}_{0}$ is defined as:
@@
Visualization. Fig. 4 (a) shows a better/worse data pair sampled from the MotionPercept dataset. The motion on the left (annotated as ‘better’ and visually superior) exhibits physics issues (e.g. floating, skating) in the simulator, while the motion on the right (annotated as ‘worse’ and visually inferior) demonstrates greater physical plausibility when simulated. Fig. 4 (b) shows three MotionPercept samples in a group that are all annotated as ‘worse’, but reveal different physical characteristics in the simulator. Our PP-Motion scores successfully capture these physical distinctions. Note that for our PP-Motion score, the higher reveals the better.
@@ Appendix A Implementation Details
We train our PP-Motion model using the MDM subset of MotionPercept, which contains 46761 better-worse motion pairs. The MDM dataset is generated from 12 action labels in HumanAct12 and 40 labels in UESTC. We first categorize the dataset according to 52 prompt labels. During training, each batch contains motions exclusively from the same category, and the PLCC loss is computed collectively across all motions within a batch. Our model architecture follows (Wang et al., 2025), employing a DSTformer (Zhu et al., 2023b) backbone with 3 layers and 8 attention heads for generating motion embedding. The embeddings are then fed into a 1024-channel MLP layer that produces a single scalar score. We train the PP-Motion model for 200 epochs with a batch size of 64 and a learning rate initialized at 4e-5 and decayed exponentially (factor 0.995 per epoch). The correlation loss term is weighted by $\lambda=0.3$.
@@ D.2. Visualize cases
We provide visualized cases from the MotionPercept dataset in Fig. 4 to verify the effectiveness of our PP-Motion metric.
@@
Our PP-Motion scoring effectively captures these physical plausibility considerations:
@@
PP-Motion scoring demonstrates strong correlation with physical feasibility assessments: